ROVIBRATIONAL DYNAMICS OF ORTHOHYDROGEN-WATER COMPLEXES IN SOLID PARAHYDRO-GEN

MARIO E. FAJARDO, AFRL/MNME, Energetic Materials Branch, Ordnance Division, U.S. Air Force Research Lab, 2306 Perimeter Rd., Eglin AFB, FL 32542-5910. mario.fajardo@eglin.af.mil.

Analysis of the rovibrational infrared (IR) absorption spectra of water (H_{2}O, D_{2}O) molecules isolated in solid parahydrogen (pH_{2}) reveals their existence as very slightly hindered rotors, with rotational constants reduced by only 2-5% from their gas phase values. Clustering of residual orthohydrogen (oH_{2}) molecules with water monomers results in the appearance of several new IR absorption features. For Type B water monomer bands (e.g. ν_{1}, ν_{2}) most of the new features appear near the vibrational origin, and were originally interpreted as indicating the presence of ”non-rotating” water molecules.^{a} However, for Type A bands (e.g. ν_{3}, ν_{2} + ν_{3}, 2ν_{2} + ν_{3}, ν_{1} + ν_{3}) very little IR activity is observed near the vibrational origin, refuting this appealingly simple explanation. Here we propose a new interpretation which assumes a semi-rigid C_{2v} structure for the ground state of the oH_{2}-water complexes, with the oH_{2} acting as a proton donor to the water oxygen atom. In this picture, the oH_{2}-water complex spectra can be understood as parallel and perpendicular bands of an asymmetric top near the prolate symmetric top limit.^{b} Thus, the features bunching near the Type B vibrational band origins arise from the ΔK = 0 selection rule for parallel bands, while the more widely separated features in the Type A bands arise from the ΔK = ±1 selection rule for perpendicular bands.

aM.E. Fajardo, 57th International Symposium on Molecular Spectroscopy, paper MF01 (2002).

bM.E. Fajardo, S. Tam, and M.E. DeRose, J. Mol. Struct. accepted (2004).